OpenStudy (bananas1234):

the length of a pendulum varies directly with the square of the period of the pendulum. If a pendulum 2 m long has a period of 4 seconds (s) what is the length of the petroleum with a peroid of 10 s

OpenStudy (bananas1234):

@pooja195 @zepdrix

OpenStudy (bananas1234):

@sshayer

OpenStudy (sshayer):

Let t be the time period and l be the length of pendulum.Then l=kt^2 2=k*4^2 \[k=\frac{ 2 }{ 4^2 }=\frac{ 2 }{ 16 }=\frac{ 1 }{ 8 }\] \[l=\frac{ 1 }{ 8 }t^2\] put t=10 and find l

OpenStudy (bananas1234):

12.5? @sshayer

OpenStudy (sshayer):

correct. l=12.5 m